A rectangular piece of metal has an area of 18 square centimeters. Its perimeter is 18 centimeters. What are the dimensions of the piece?
Answer by edjones(8007)

L*w=18
.
2L+2w=18
2L=18-2w
L=9-w
.
w(9-w)=18
9w-w^2=18
w^2-9w+18=0
(w-6)(w-3)=0
L=6 and w=3 or vice versa.
.
Ed